# Smooth polynomials with several prescribed coefficients

### v3

@@ 5.1. Major arc estimates
$\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(\frac{af}{g}\right)$ $=\sum_{d\mid g}\frac{\mu(g/d)}{q^{\ell}\Phi(g/d)}\Psi_{g/d}(n-\deg d,m)+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n}e^{\varepsilon(\ell+\deg g)}|g|^{1/2}\right),$
@@
where in the innermost sum $c\in{\mathcal{M}}(n-\deg d)$ runs over all resuced residue calsses modulo ${\mathcal{R}}_{g/d,\ell}$ such that $c\equiv b\mod{\mathcal{R}}_{\ell}$. As $n-\deg d>\deg(g/d)+\ell$, by (19), the polynomial $c$ runs thorough all residue classes modulo $g/d$. Thus by Lemma 11, the contributions of the principal characters $\chi_{0}^{(g/d,\ell)}$ modulo ${\mathcal{R}}_{g/d,\ell}$ is
@@
By (19) and Lemma 18, the contribution of the non-principal characters are
@@
By the Cauchy-Schwartz inequality and Lemma 12 we have
@@
$\ll q^{\left(\frac{1}{2}+\varepsilon\right)n}e^{\varepsilon(\ell+\deg g)}\sum_{d\mid g}q^{-\left(\frac{1}{2}+\varepsilon\right)\deg d}\Phi(g/d)^{1/2}\ll q^{\left(\frac{1}{2}+\varepsilon\right)n}e^{\varepsilon(\ell+\deg g)}|g|^{1/2}\sum_{d\mid g}\frac{1}{|d|^{1+\varepsilon}}$
@@
$+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n+k/2}e^{\varepsilon(\ell+k}\right),$
@@ 5.2. Minor arc estimates
$\left\{|e|=\frac{q^{i}}{|g|}:\|eg\gamma+\beta\|<q^{-j}\right\}\ll\max\left\{1,\frac{q^{i-j}}{|g|},\|g\gamma\|\frac{q^{i}}{|g|},\frac{1}{\|g\gamma\|q^{j}}\right\}$
@@
$S(\xi;n,m)\ll m^{2}n^{1/2}\left(|g|^{1/2}q^{3n/4+m/4+1/2}+\|g\gamma\|^{1/2}q^{n}+\frac{q^{(n+1)/2}}{\|\gamma\|^{1/2}}\right)$
@@ 6. Analysis of $S_{\mathcal{J}}$
We have the following results, for their proof see the proof of [12, Section 2.1].
@@
The following result is [12, Lemma 2.7]
@@
Let $a,g\in\mathbb{F}_{q}[t]$ be two polynomials with $\gcd(a,g)=1$ and write $g=g_{0}t^{k}$, $\gcd(g_{0},t)=1$. If $1<|g_{0}|\leq q^{\lceil n/(\#{\mathcal{I}}+1)\rceil}-1$, then
@@
The following result is the analogue of Bourgain’s result [2, Lemma 3] proved by Ha [12, Lemma 2.6]
@@
$\sum_{\begin{subarray}{c}|g|<q^{k}\\ \gcd(aT,g)=1\end{subarray}}\left|S_{\mathcal{J}}\left(\frac{a}{g}\right)\right|\leq\frac{q^{n}}{q^{\#{\mathcal{I}}}}q^{2kC_{1}\#{\mathcal{I}}/n},$
@@
$\kappa(x,y,\delta)=\left\{\begin{array}[]{cl}\frac{2}{\delta+1}&\text{if }1<x/y<2\\ \frac{2v}{(v+1)\delta+v-1}&\text{if }v-1<x/y<v\\ 1&\text{if }y\mid x.\end{array}\right$.
@@ 7. Proof of Theorem 4
As the polynomials $f$ in the summations $S(\xi;n,m)$ and $S_{\mathcal{J}}(\xi)$ are monic and of degree $n$, we have for $\xi\in{\mathfrak{M}}$ having the form $\xi=a/g+\gamma$, $|\gamma|<q^{-n+\ell}$ that $\mathbf{e}(f\gamma)=\mathbf{e}(b\gamma)$ if $f\equiv b\mod{\mathcal{R}}_{\ell}$, $b\in{\mathcal{M}}(n)$. Then
@@
(35) $\begin{split}M_{1,0}=&\frac{1}{q^{n-\ell}}\sum_{\begin{subarray}{c}b\bmod{\mathcal{R}}_{\ell}\\ b\in{\mathcal{M}}(n)\end{subarray}}\left(\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}1\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}1\right)\\ =&\frac{1}{q^{n-\ell}}\left(\frac{1}{q^{\ell}}\Psi(n,m)+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n}e^{\varepsilon\ell}\right)\right)\sum_{\begin{subarray}{c}b\bmod{\mathcal{R}}_{\ell}\\ b\in{\mathcal{M}}(n)\end{subarray}}\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}1\\ =&\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n-\#{\mathcal{I}}+\ell}e^{\varepsilon\ell}\right).\end{split}$
Now consider the terms with $M_{1,k}$ with $k\geq 1$. Let $B={\mathcal{I}}\cap\{n-\ell,\dots,n-1\}$. Clearly, one can represent the residue classes of $\mathcal{R}$$_{\ell}$ by the elements of ${\mathcal{M}}(\ell)$. The contribution of those $b$ that $b_{i-(n-\ell)}\neq\alpha_{i}$ for $i\in B$ is zero. Write
@@
$E(n,m,k)\ll q^{\left(\frac{1}{2}+\varepsilon\right)n+k/2}e^{\varepsilon(\ell+k)}.$
@@
(41) $\begin{split}M_{1}=&\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n-\#{\mathcal{I}}+\ell}e^{\varepsilon\ell}\right)+\sum_{\begin{subarray}{c}1\leq k\leq\kappa\\ k-1\in{\mathcal{I}}_{\ell}\end{subarray}}M_{1,k}\\ =&\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n-\#{\mathcal{I}}+\ell}e^{\varepsilon\ell}\right)\\ &+(q-1)\sum_{\begin{subarray}{c}0\leq j<\nu\\ i_{j}<\kappa\end{subarray}}\frac{1}{q^{\#{\mathcal{I}}-j-\ell}}\Bigg{(}\frac{1}{q^{\ell-1}(q-1)}\Psi(n-i_{j}-1,m)-\frac{1}{q^{\ell}(q-1)}\Psi(n-i_{j},m)\\ &\qquad\qquad\qquad\qquad\qquad\qquad+E(n,m,i_{j}+1)\Bigg{)}\\ &-\mathbf{1}(i_{\nu}<\kappa)\frac{1}{q^{\#{\mathcal{I}}-\nu-\ell}}\Bigg{(}\frac{1}{q^{\ell-1}(q-1)}\Psi(n-i_{\nu}-1,m)-\frac{1}{q^{\ell}(q-1)}\Psi(n-i_{\nu},m)\\ &\qquad\qquad\qquad\qquad\qquad\qquad+E(n,m,i_{\nu}+1)\Bigg{)}\\ =&\frac{1}{q^{\#{\mathcal{I}}}}\Psi(n,m)+\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\Lambda(n,m,{\mathcal{I}})+O\left(q^{\left(-\frac{1}{2}+\varepsilon\right)n+\kappa/2+\ell+\nu}e^{\varepsilon(\ell+\kappa)}\right)\right).\end{split}$
For $M_{2}$, write $K=\lceil n/(\#{\mathcal{I}}+1)\rceil$. If $g$ has the form $g=g_{0}t^{k}$ with $t\nmid g_{0}$ and $1<|g_{0}|<q^{K}$, then $S_{{\mathcal{J}}}(a/g)$ vanishes by Lemma 26. Then, we can write
(42) $M_{2}=\frac{1}{q^{n-\ell}}\sum_{r=K}^{\kappa}\sum_{k=0}^{\kappa-r}M_{2}(r,k)$
@@
(43) $R(r,k)\ll q^{\left(\frac{1}{2}+\varepsilon\right)n+\frac{r+k}{2}}e^{\varepsilon(\ell+\kappa)}.$
By (4), (20) and Lemma 14, we have
@@
Clear, (7) is zero, unless,
@@
(50) $\begin{split}M_{2}&\ll\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\rho\left(\frac{n}{m}-\frac{n}{2m^{2}}\right)e^{T(n,m)}\sum_{r\geq K}\frac{r\left(1+\log_{q}r\right)}{q^{C_{2}r}}+O\left(\kappa q^{\left(-\frac{1}{2}+\varepsilon\right)n+\kappa+\frac{1}{2}+\ell}e^{\varepsilon(\ell+\kappa)}\right)\right)\\ &\ll\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\rho\left(\frac{n}{m}-\frac{n}{2m^{2}}\right)e^{T(n,m)}\frac{K\left(1+\log_{q}K\right)}{q^{C_{2}K}}+O\left(\kappa q^{\left(-\frac{1}{2}+\varepsilon\right)n+\kappa+\frac{1}{2}+\ell}e^{\varepsilon(\ell+\kappa)}\right)\right).\end{split}$
@@
(51) $\begin{split}\int_{{\mathfrak{M}}}S(\xi;n,m)\overline{S_{\mathcal{J}}(\xi)}\mathrm{d}\xi&=\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}\left(1+O\left(e^{T(n,m)}\frac{\delta^{-1}\log\delta}{q^{C/\delta}}\right)\right)\\ &\quad+\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\Lambda(n,m,{\mathcal{I}})+O\left(\kappa q^{\left(-\frac{1}{2}+\varepsilon\right)n+\kappa+\ell+\nu+\frac{1}{2}}e^{\varepsilon(\ell+\kappa)}\right)\right).\end{split}$
@@
(53) $O\left(\kappa q^{\left(-\frac{1}{2}+\varepsilon\right)n+\kappa+\ell+\nu+\frac{1}{2}}e^{\varepsilon(\ell+\kappa)}\right)=O\left(\kappa q^{\left(-\frac{1}{2}+\varepsilon\right)n+3\kappa+\frac{3}{2}}e^{2\varepsilon\kappa}.\right)$
@@
If $\xi\in{\mathfrak{m}}$, then either $q^{\kappa}\leq|g|<q^{n/2}$ or $|g|<q^{\kappa}$ and $q^{-n+\ell}\leq|\gamma|\leq 1/(|g|q^{n/2})$. In the formal case, we estimate $S(\xi;n,m)$ by the first bound of Lemma 23

### v4

@@ 5.1. Major arc estimates
$\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(\frac{af}{g}\right)$ $=\sum_{d\mid g}\frac{\mu(g/d)}{q^{\ell}\Phi(g/d)}\Psi_{g/d}(n-\deg d,m)+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n}|g|^{1/2}\right),$
@@
where in the innermost sum $c\in{\mathcal{M}}(n-\deg d)$ runs over all reduced residue classes modulo ${\mathcal{R}}_{g/d,\ell}$ such that $c\equiv b\mod{\mathcal{R}}_{\ell}$. As $n-\deg d>\deg(g/d)+\ell$, by (21), the polynomial $c$ runs through all all residue classes modulo $g/d$. Thus by Lemma 11, the contributions of the principal characters $\chi_{0}^{(g/d,\ell)}$ modulo ${\mathcal{R}}_{g/d,\ell}$ is
@@
By (21) and Corollary 18, the contribution of the non-principal characters are
@@
By the Cauchy-Schwarz inequality and Lemma 12 we have
@@
$\ll q^{\left(\frac{1}{2}+\varepsilon\right)n}\sum_{d\mid g}q^{-\left(\frac{1}{2}+\varepsilon\right)\deg d}\Phi(g/d)^{1/2}\ll q^{\left(\frac{1}{2}+\varepsilon\right)n}|g|^{1/2}\sum_{d\mid g}\frac{1}{|d|^{1+\varepsilon}}$
@@
$+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n+k/2}\right),$
@@ 5.2. Minor arc estimates
$\left\{e:\deg e=i-\deg g,\ \|eg\gamma+\beta\|<q^{-j}\right\}\ll\max\left\{1,\frac{q^{i-j}}{|g|},\|g\gamma\|\frac{q^{i}}{|g|},\frac{1}{\|g\gamma\|q^{j}}\right\}$
@@
$S(\xi;n,m)\ll m^{2}n^{1/2}\left(|g|^{1/2}q^{3n/4+m/4+1/2}+\|g\gamma\|^{1/2}q^{n}+\frac{q^{(n+1)/2}}{\|\gamma\|^{1/2}}\right).$
@@ 6. Analysis of $S_{\mathcal{J}}$
We have the following results, for their proof see the proof of [18, Section 2.1].
@@
The following result is [18, Lemma 2.7].
@@
Let $a,g\in\mathbb{F}_{q}[t]$ be two polynomials with $\gcd(a,g)=1$ and write $g=g_{0}t^{k}$, $\gcd(g_{0},t)=1$. If $1<|g_{0}|\leq q^{\lceil(n-k)/(\#{\mathcal{I}}+1)\rceil}-1$, then
@@
The original result [18, Lemma 2.7] is stated correctly for $k=0$; however, for $k\geq 1$ it requires a minor revision.
Proof.
Write
$\frac{a}{g}=\sum_{i\geq 0}\xi_{-i}t^{-i}.$
Assume, that $S_{\mathcal{J}}(a/g)\neq 0$. Then, by Lemma 25, we have
$\xi_{-i-1}=0\quad\text{for }i\in\{0,\dots,n-1\}\setminus{\mathcal{I}}$
In addition, as $g=g_{0}t^{k}$, we also have
$\xi_{0}=\dots=\xi_{k-1}=0.$
Put $J=\lceil(n-k)/(\#{\mathcal{I}}+1)\rceil$. By the pigeon-hole principle, there is at least $\lceil(n-\#{\mathcal{I}}-k)/(\#{\mathcal{I}}+1)\rceil\geq J-1$ consecutive indices in the range $\{k,\dots n\}$ where the coefficients $\xi_{i}$ vanish, say
$\xi_{r+1}=\dots=\xi_{r+J-1}=0.$
Then
$\left|\left\{t^{r}\cdot\frac{a}{g}\right\}\right|\leq\left|\frac{1}{t^{J}}\right|\leq q^{-J},$
where $\{x\}$ denotes the fractional part of $x$. If $|g_{0}|\neq 1$, the left hand side is at least $|g_{0}|^{-1}$ whence $|g_{0}|\geq q^{J}$. ∎
The following result is the analogue of Bourgain’s result [2, Lemma 3] proved by Ha [18, Lemma 2.6]
@@
$\sum_{\begin{subarray}{c}|a|<|g|<q^{k}\\ \gcd(aT,g)=1\end{subarray}}\left|S_{\mathcal{J}}\left(\frac{a}{g}\right)\right|\leq\frac{q^{n}}{q^{\#{\mathcal{I}}}}q^{2kC_{1}\#{\mathcal{I}}/n},$
@@
$\kappa(x,y,\delta)=\left\{\begin{array}[]{cl}\frac{2}{\delta+1}&\text{if }1<x/y<2\\ \frac{2v}{(v+1)\delta+v-1}&\text{if }v-1<x/y<v\text{ for some integer $v$}\\ 1&\text{if }y\mid x.\end{array}\right$.
@@ 7. Proof of Theorem 4
As the polynomials $f$ in the summations $S(\xi;n,m)$ and $S_{\mathcal{J}}(\xi)$ are monic and of degree $n$, we have for $\xi\in{\mathfrak{M}}$ having the form $\xi=a/g+\gamma$, $|\gamma|<q^{-n+\ell}$ that $\mathbf{e}(f\gamma)=\mathbf{e}(b\gamma)$ if $f\equiv b\mod{\mathcal{R}}_{\ell}$, $b\in{\mathcal{M}}(n)$ by the choice (19) of $\ell$. Then
@@
(38) $\begin{split}M_{1,0}=&\frac{1}{q^{n-\ell}}\sum_{\begin{subarray}{c}b\bmod{\mathcal{R}}_{\ell}\\ b\in{\mathcal{M}}(n)\end{subarray}}\left(\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}1\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}1\right)\\ =&\frac{1}{q^{n-\ell}}\left(\frac{1}{q^{\ell}}\Psi(n,m)+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n}\right)\right)\sum_{\begin{subarray}{c}b\bmod{\mathcal{R}}_{\ell}\\ b\in{\mathcal{M}}(n)\end{subarray}}\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}1\\ =&\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n-\#{\mathcal{I}}+\ell}\right).\end{split}$
Now consider the terms with $M_{1,k}$ with $k\geq 1$. Let $B={\mathcal{I}}\cap\{n-\ell,\dots,n-1\}$. Clearly, one can represent the residue classes of $\mathcal{R}$$_{\ell}$ by the elements of ${\mathcal{M}}(\ell)$. The contribution of those $b$ such that $b_{i-(n-\ell)}\neq\alpha_{i}$ for $i\in B$ is zero. Write
@@
$E(n,m,k)\ll q^{\left(\frac{1}{2}+\varepsilon\right)n+k/2}.$
@@
(44) $\begin{split}M_{1}=&\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n-\#{\mathcal{I}}+\ell}\right)+\sum_{\begin{subarray}{c}1\leq k\leq\kappa\\ k-1\in{\mathcal{I}}_{\ell}\end{subarray}}M_{1,k}\\ =&\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n-\#{\mathcal{I}}+\ell}\right)\\ &+(q-1)\sum_{\begin{subarray}{c}0\leq j<\nu\\ i_{j}<\kappa\end{subarray}}\frac{1}{q^{\#{\mathcal{I}}-j-\ell}}\Bigg(\frac{1}{q^{\ell-1}(q-1)}\Psi(n-i_{j}-1,m)-\frac{1}{q^{\ell}(q-1)}\Psi(n-i_{j},m)\\ &\qquad\qquad\qquad\qquad\qquad\qquad+E(n,m,i_{j}+1)\Bigg)\\ &-\mathbf{1}(i_{\nu}<\kappa)\frac{1}{q^{\#{\mathcal{I}}-\nu-\ell}}\Bigg(\frac{1}{q^{\ell-1}(q-1)}\Psi(n-i_{\nu}-1,m)-\frac{1}{q^{\ell}(q-1)}\Psi(n-i_{\nu},m)\\ &\qquad\qquad\qquad\qquad\qquad\qquad+E(n,m,i_{\nu}+1)\Bigg)\\ =&\frac{1}{q^{\#{\mathcal{I}}}}\Psi(n,m)+\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\Lambda(n,m,{\mathcal{I}})+O\left(q^{\left(-\frac{1}{2}+\varepsilon\right)n+\kappa/2+\ell+\nu}\right)\right).\end{split}$
Next, we consider $M_{2}$. For the summation $g$, write $g=g_{0}t^{k}$ with $t\nmid g_{0}$. By Lemma 26, if $g_{0}<q^{\lceil(n-k)/(\#{\mathcal{I}}+1)\rceil}$, $S_{{\mathcal{J}}}(a/g)$ vanishes. Put
(45) $R_{k}=\left\lceil\frac{n-k}{\#{\mathcal{I}}+1}\right\rceil$
Then, we can write
(46) $M_{2}=\frac{1}{q^{n-\ell}}\sum_{k=0}^{\kappa}\sum_{r=R_{k}}^{\kappa-k}M_{2}(r,k)$
@@
(47) $R(r,k)\ll q^{\left(\frac{1}{2}+\varepsilon\right)n+\frac{r+k}{2}}$
Since $u\geq 100$ by (13), we have by (4), (22) and Lemma 14 that
@@
Clearly, (7) is zero, unless,
@@
(54) $\begin{split}M_{2}&\ll\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\rho\left(\frac{n}{m}-\frac{n}{2m^{2}}\right)e^{T(n,m)}\sum_{r\geq R_{1}}\frac{r\left(1+\log_{q}r\right)}{q^{C_{2}r}}+O\left(\kappa q^{\left(-\frac{1}{2}+\varepsilon\right)n+\kappa+\frac{1}{2}+\ell}\right)\right)\\ &\ll\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\rho\left(\frac{n}{m}-\frac{n}{2m^{2}}\right)e^{T(n,m)}\frac{R_{1}\left(1+\log_{q}R_{1}\right)}{q^{C_{2}R_{1}}}+O\left(\kappa q^{\left(-\frac{1}{2}+\varepsilon\right)n+\kappa+\frac{1}{2}+\ell}\right)\right),\end{split}$
where $R_{1}$ is defined by (45).
@@
(55) $\begin{split}\int_{{\mathfrak{M}}}S(\xi;n,m)\overline{S_{\mathcal{J}}(\xi)}\mathrm{d}\xi&=\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}\left(1+O\left(e^{T(n,m)}\frac{\delta^{-1}\log\delta}{q^{C/\delta}}\right)\right)\\ &\quad+\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\Lambda(n,m,{\mathcal{I}})+O\left(\kappa q^{\left(-\frac{1}{2}+\varepsilon\right)n+\kappa+\ell+\nu+\frac{1}{2}}\right)\right)\end{split}$
for some positive $C$.
@@
(57) $O\left(\kappa q^{\left(-\frac{1}{2}+\varepsilon\right)n+\kappa+\ell+\nu+\frac{1}{2}}e^{\varepsilon(\ell+\kappa)}\right)=O\left(\kappa q^{\left(-\frac{1}{2}+\varepsilon\right)n+3\kappa+\frac{3}{2}}\right).$
@@
If $\xi\in{\mathfrak{m}}$, then either $q^{\kappa}\leq|g|<q^{n/2}$ or $|g|<q^{\kappa}$ and $q^{-n+\ell}\leq|\gamma|\leq 1/(|g|q^{n/2})$. In the former case, we estimate $S(\xi;n,m)$ by the first bound of Lemma 23